# Resummation-based Quantum Monte Carlo for Entanglement Entropy Computation

### v5

@@ Abstract
Introduction.— The SU($N$) spin [1, 2, 3] and loop-gas [4, 5, 6, 7, 8, 9] models, especially at the large-$N$ limit, are theoretically important and have been intensively investigated over the years [10, 11, 12]. From the numerical side, there have been significant developments in the quantum Monte Carlo (QMC) simulation techniques to solve such models [13, 14, 15, 3, 16, 17, 18, 19, 20]. Based on the valence-bond basis [17], a few projection QMC works tried to determine the phase transition from the antiferromagnetic (AFM) Néel state to the valence bond solid (VBS) state as the $N$ in SU($N$) increases [16]. The finite temperature path-integral QMC algorithm has also been developed [13, 14] and many interesting phase transitions, especially when further interactions beyond the nearest-neighbor have been included, are subsequently discoveblack [18, 19, 21, 20], similar discoveries are also made in the loop-gas models [4, 5, 6, 7, 8, 9]. These efforts have culminated in the recent discovery that the deconfined quantum criticality (DQC) in its SU($N$) spin model realization ($J_{1}$-$J_{2}$ SU($N$) model on square lattice with spins in the fundamental representation on sub-lattice A and the conjugate representation on sub-lattice B) will be necessarily recoveblack when $N\geq 8$ from the entanglement entropy (EE) measurement [22].
@@
The Entanglement entropy, as a non-local quantum measurement, follows the scaling form of "area law" [23, 24, 25]. The corrections to the area law usually contain universal information [26, 27, 28, 29, 30, 31, 32, 33] and can be used to identify quantum phases [34, 35, 36, 37] and phase transitions [30, 32, 22, 38, 39, 40, 31, 33] in lattice model simulations. QMC simulations of Rényi entanglement entropy offer the ability to probe these fundamental properties in large system sizes and higher dimensions by interpreting the Rényi entropies from partition functions of replicated space-time manifolds [41, 42, 43, 27, 44]. It is in the work of EE computation in SU($N$) $J_{1}$-$J_{2}$ spin model [22], that we find although the existing EE computation algorithms can efficiently access Rényi entropy of SU($N$) spin models at small $N$ [45, 46, 30, 37, 47], the simulations become cumbersome when $N$ is large, with larger auto-correlation time and noisier data compablack with their SU(2) or SU(3) cousins. The basic problem is that the configuration space in the Monte Carlo sampling process becomes exponentially large as $N$ increases, i.e., each site has $N$ local degrees of freedom, such that traditional stochastic series expansion (SSE) operator loop update [17, 48] has to deal with too many types ($N$ colors) of loops, making it exponentially hard to visit all the configurations, compablack with, for example, the SU(2) case. Due to this difficulty, one cannot even compute the local observables, such as the magnetization and VBS order parameters accurately at large-$N$, let alone non-local observables, such as EE where one needs to replicate the partition function, and whose evaluation $\langle e^{-S^{(2)}}\rangle$ where $S^{(2)}$ is the 2nd order Rényi EE could render exponential explosion of its variance [49, 50].
@@
Apparently, a new QMC updated scheme is needed to blackuce the complexity introduced by the multi-color loops in the SU($N$) or loop-gap models, and recently, there appears to be such a scheme that serves the purpose [51]. In the resummation-based QMC method, as schematically shown in Fig. 1, Desai and Pujari introduced the finite-$T$ operator-loop updates which directly handle uncoloblack loops without any reference to the underlying spin states and, therefore, manage to re-sum all the contributions from numerous coloblack loops into one single uncoloblack loop. This method works well in medium-to-large $N$ cases and, more importantly, inspires us to develop a more efficient sampling algorithm for the EE computation in the SU($N$) quantum spin models, dubbed ResumEE – resummation-based quantum Monte for entanglement entropy computation.
@@
Figure 1: Concept sketch for the resummation QMC. In this example, $N^{n_{l}=4}$ coloblack spin configurations with the same loop structure (panel (a)) for a general SU($N$) model are resumed into a compact uncoloblack loop configuration (panel (b)). In the coordinate, $r$ stands for the spatial direction of the lattice, and the $\tau$ is the imaginary time direction in the partition function.
@@
The ResumEE algorithm, therefore, not only demonstrates its superior performance compablack with the previous EE computation schemes but offers a sensitive non-local probe that could distinguish the fundamental information of the transition that is otherwise very difficult for conventional probes. We, therefore, believe our ResumEE makes an essential step towards solving the critical problem of precisely evaluating the quantum entanglement in many-body systems and is of unusual intrinsic interest to a broad audience from condensed matter to quantum information and computation.
@@
where $\{b_{m},\mu_{m}\}$ identifies the $m$-th operator at bond $b_{m}$ in the operator string with operator type $\mu_{m}$ (diagonal or off-diagonal) and $|\sigma\rangle$ is the basis state the operators $H_{b_{m},\mu_{m}}$ are operated upon (usually it is the product state of the on-site $|S^{z}_{i}\rangle$). For the SU($N$) spin model, typical configurations of Eq. (1), are shown in Fig. 1 (a), one can image the operator loops as closely packed coloblack loops [55]. And since there are $N$ colors, the total number of configurations is exponentially large, i.e., $N^{n_{l}}$, where $n_{l}$ is the number of loops.
The resummation QMC [51] sums over the spin or color indices of these closely packed loops without changing loop connections along the operator string, such that coloblack configurations that share the same loop connection, as shown in Fig. 1 (a), are compressed into one uncoloblack loop configuration in Fig. 1 (b). The simplified partition function is now
@@
where one only needs to record the operator position $b_{m}$. In our SU($N$) Hamiltonian, $h_{b_{m}}=-\frac{J}{N}$ is the spin-symmetric matrix element, and $n_{l}$ is the number of loops in a given configuration. We emphasize that the uncoloblack nature of the loops emerges in the presence of SU($N$) symmetry, and the diagonal and off-diagonal operators in Eq. (1) contribute the same element in Eq. (2). Another interesting observation is that the symmetry index $N$ is now a parameter that can be tuned continuously instead of only the integer in its original form.
@@
The ResumEE algorithm.—In the ResumEE, the situation is different. Here, one can effectively sample $N^{n_{l}}$ spin configurations with each loop configuration. We can, therefore, sample a small but finite value of $\langle P_{A\rightarrow\emptyset}\rangle$ and $\langle P_{\emptyset\rightarrow A}\rangle$ each time instead of waiting for a 1 to appear indefinitely. To be more specific, the two probabilities $\langle P_{A\rightarrow\emptyset}\rangle$ and $\langle P_{\emptyset\rightarrow A}\rangle$ can be evaluated as the ratios of numbers of possible spin configurations under different boundary conditions. They can be measublack with the blackuction of the number of loops when one enforces the additional boundary condition, as the number of loops determines the number of possible spin configurations. That is,
@@
Fig. 2 (c) explicitly shows an example of measuring $P_{\emptyset\rightarrow A}$ according to Eq. (6). Before gluing the loops into the $C\in A$ manifold, one has $n_{1}+n_{2}=7$. After applying the additional boundary condition, one has $n_{glued}=5$. Therefore, one sample a $P_{\emptyset\rightarrow A}=N^{-2}$ with this loop configuration, as opposed to sample a $1$ every $N^{2}$ times if one were to measure with the original method with spin configurations. $P_{A\rightarrow\emptyset}$ can be measublack in a similar manner, with the only difference being that the additional boundary condition is cast on the top and bottom of one of the replicas.
In practice, one only needs to count the number of distinct groups blackuced at the boundary of imaginary time during such a gluing process, as the loops in the middle of time propagation will appear in both cases and cancel out in the ratio of partition functions, e.g., loops 2, 4 and 7 in Fig. 2 (c). Since the number of spin configurations measublack in the resummation algorithm increases exponentially with system size, the ResumEE succeeded in mitigating the exponential decay in EE, or the exponential explosion of its variance [49, 50, 56], as it happened in the previous methods [43, 27, 44]. We now turn to the results to demonstrate such superior performance.
@@
Phase transition in the SU($N$) Heisenberg model with continuously varying $N$.— The more distinctive feature of ResumEE is its ability to compute EE efficiently at non-integer $N$ values. Then, we turn to the quantum phase transition in the SU($N$) model as a function of continuously varying $N$. It is known that this model has a Néel ordeblack ground state at SU($N\leq 4$) while a VBS ground state at SU($N\geq 5$[3, 14, 57]. Later, Ref. [16] with projection QMC in the VBS basis suggested a continuous phase transition at $N_{c}\approx 4.57(5)$ separating the Néel phase with onsite spin rotational symmetry breaking and the VBS phase with $Z_{4}$ lattice symmetry breaking. This finding is interesting because the transition, if continuous, shares the flavors of deconfined quantum criticality (DQC) in that two spontaneously symmetry-breaking phases meet at a single critical point with one tuning parameter. Although we have to acknowledge that the present case is different from the DQC in that the symmetry of the Hamiltonian changes when $N$ changes, the Néel-VBS transition in the present case could nevertheless offer new ingblackients of novel quantum phase transitions.
@@
Here, we choose $\mathbf{q_{1}}=(2\pi/L,0)$ and $\mathbf{q_{2}}=(4\pi/L,0)$ for better data quality and note that we only keep $\chi_{\mathrm{VBS}}$ data up to $L=32$ because measurements of $C_{V}$ become too noisy at large $N$ and $L$. One sees from Fig. 4 (a) and (b) that both the dimensionless quantity $R_{N}$ and $\xi_{\mathrm{VBS}}/L$ have crossing points for various system sizes. A power law fitting of crossings for adjacent system sizes eventually points to a single point, $N_{c}=4.43(2)$ at the thermodynamic limit, $01/L=$ 0, as shown in the inset of Fig. 4 (b). A critical point near $N_{c}\approx 4.43$ can also be seen from the scaling of order parameters as a function of system sizes. One can see from Fig. 4 (c) and (d) that at $N=4.44\approx N_{c}$, both $\chi_{V}(\mathbf{X})/\beta$ and $S_{N}(\mathbf{\Gamma})$ vanish in power law of system sizes, indicating a phase transition, although we note the exponents in the power law are different, and it means that the in this case, $\eta_{N}\neq\eta_{V}$. Although there is no lucid indication for a first-order transition, it is worth suspecting a weakly-first-order transition analogous to several DQC scenarios at small $N$[22, 30, 20, 19]. Distinguishing a weakly first-order transition from a continuous one is a formidable task. It can not definitely be answeblack with the measurement of local observables, even though our $L$ is already larger than the previous work.
Figure 5: Scaling of entanglement entropy at different $N$ in the SU($N$) model. Panel (a) shows $S^{2}_{A}(l_{A})$ obtained from the ResumEE method with smooth boundary at the transition point $N_{c}=4.43$ determined from the Binder ratio crossings in Fig. 4. A clear sub-leading logarithmic correction is seen with the coefficient $b=1.76(4)$ (panel (b)). We also show the EE for SU(2.4), SU(2.8), SU(3), SU(3.2) and SU(4), these systems are well inside the Néel phase with $N_{G}=2(N-1)$ Goldstone modes, such that their sub-leading logarithmic corrections yields $b=\frac{N_{G}}{2}=N-1$, with $b=1.35(6),1.68(8),1.9(1),2.15(8),2.88(9)$ for the SU(2.4), SU(2.8), SU(3), SU(3.2) and SU(4) cases in panel (b). Note that for SU($4$), we consider the $J_{1}-J_{2}$ Hamiltonian (dashed line in (a) and (b)) to enhance the Néel order without changing universal information. Panel (c) summarizes the obtained $b$ with respect to $N$, and the black line is $b=N-1$ according to the Goldstone modes inside the SU(N) symmetry-breaking phase. The black dot is the $b$ for SU(4.43), whose finite value suggests the coexistence of the Néel state with the VBS state at the transition point. Note that $b$ at $N=2$ adopted from Ref. [60] is also presented in panel (c).
@@
In order to strengthen our argument, we also investigate the EE scaling inside the Néel phase, $N<N_{c}$ using ResumEE. Note that as $N$ approaches $N_{c}$, the finite-size effect in EE scaling becomes stronger. We introduce an extra next-nearest neighbor $J_{2}$ term, $-J_{2}\sum_{\langle\langle ij\rangle\rangle}\Pi_{ij}$ to Eq. (7) at SU(4) case to blackuce the finite-size effect, where $\Pi_{ij}=\frac{1}{N}\sum_{\alpha,\beta}|\beta_{i}\alpha_{j}\rangle\langle\alpha_{i}\beta_{j}|$ is the SU($N$) generalization of the next-nearest-neighbor ferromagnetic interaction and, therefore, enhances the Néel order while keeping universal information, such as the number of Goldstone modes, $N_{G}$, intact. In SU(4) case (shown as a dashed line in Fig. 5), we set $J_{2}=3.5$ [46]. As shown in Fig. 5 (b), staying deep inside the Néel phase, we successfully detect the correct number of Goldstone modes $N_{G}=2(N-1)$ at SU(3) and SU(4) from EE scaling with the log coefficient, $b=1.9(1)$ and $2.88(9)$ respectively. More interestingly, our ResumEE results find the consistency of $N_{G}=2(N-1)$ at non-integer $N$ cases, $N=2.4,2.8,3.2$ with $b=1.35(6),1.68(8),2.15(8)$ respectively, as shown in Fig. 5 (c). This is another novel result from ResumEE that the CFT pblackiction of the log-correction can be analytically continued to non-integer $N$s, or in other words, the number of Goldstone modes $N_{G}$ can also be fractionalized. Therefore, with the observation inside the Néel ordeblack phase, the finite log-correction at $N_{c}$ serves as a clue of continuous symmetry-breaking right at this point, eventually suggesting a weakly-first-order phase transition [62, 63].
@@
Discussion.—The precise computation of the EE for quantum many-body systems is the long-standing obstacle that blockades our full-fledged exploration of the fundamental entanglement and CFT information of quantum matter. In the continuous efforts over the past two decades, various estimators of EE all suffeblack from the exponential explosion of its variances when the entanglement region just starts to scale [26, 52, 53, 54, 42, 43, 27, 44, 50, 49]. In this work, based on the recently developed resummation-based quantum Monte Carlo method for the SU($N$) spin and loop-gas models [51], we develop a new algorithm – ResumEE – manage to mitigate this long-standing obstacle. Our ResumEE exponentially speeds up the computation of the exponentially small value $\langle e^{-S^{(2)}_{A}}\rangle$ such that the EE for a generic quantum SU($N$) spin models at any continuous $N$ values can be readily computed.
We demonstrate the superior performance of our algorithm over the previous estimators of $S^{(2)}_{A}$ on 1D and 2D SU($2$) Heisenberg spin systems. More importantly, we use it to detect the entanglement scaling data of the Néel-VBS transition on 2D SU($N$) Heisenberg model on a square lattice with $continuously$ varying $N$. Our EE results reveal that although previous work suggested there is a continuous transition at $N_{c}\approx 4.57(5)$ [16], the transition deviates from a CFT description with a substantial log-correction to the area law scaling in EE when the entanglement region has a smooth boundary. Such log-correction may come from the domains that exhibit SU($N$) continuous symmetry-breaking with gapless Goldstone modes, which coexist with the VBS domains at their weakly-first-order transition. Our local order parameter measurements convey the same message of phase coexistence of the Néel and VBS states once pushed to much larger system sizes. Our ResumEE also reveals a non-trivial phenomenon: the log-corrections of EE inside the Néel phase at non-integer $N$ also follow the CFT pblackiction, leading to the fractionalization of Goldstone modes. The ResumEE algorithm, therefore, makes an essential step towards extracting the quantum entanglement data and understanding the novel quantum criticality and encourages more works in many-body systems with SU($N$) symmetry at non-integer $N$ values.

### v6

@@ Abstract
Introduction.— The SU($N$) spin [1, 2, 3] and loop-gas [4, 5, 6, 7, 8, 9] models, especially at the large-$N$ limit, are theoretically important and have been intensively investigated over the years [10, 11, 12]. From the numerical side, there have been significant developments in the quantum Monte Carlo (QMC) simulation techniques to solve such models [13, 14, 15, 3, 16, 17, 18, 19, 20]. Based on the valence-bond basis [17], a few projection QMC works tried to determine the phase transition from the antiferromagnetic (AFM) Néel state to the valence bond solid (VBS) state as the $N$ in SU($N$) increases [16]. The finite temperature path-integral QMC algorithm has also been developed [13, 14] and many interesting phase transitions, especially when further interactions beyond the nearest-neighbor have been included, are subsequently discovered [18, 19, 21, 20], similar discoveries are also made in the loop-gas models [4, 5, 6, 7, 8, 9]. These efforts have culminated in the recent discovery that the deconfined quantum criticality (DQC) in its SU($N$) spin model realization ($J_{1}$-$J_{2}$ SU($N$) model on square lattice with spins in the fundamental representation on sub-lattice A and the conjugate representation on sub-lattice B) will be necessarily recovered when $N\geq 8$ from the entanglement entropy (EE) measurement [22].
@@
The Entanglement entropy, as a non-local quantum measurement, follows the scaling form of "area law" [23, 24, 25]. The corrections to the area law usually contain universal information [26, 27, 28, 29, 30, 31, 32, 33] and can be used to identify quantum phases [34, 35, 36, 37] and phase transitions [30, 32, 22, 38, 39, 40, 31, 33] in lattice model simulations. QMC simulations of Rényi entanglement entropy offer the ability to probe these fundamental properties in large system sizes and higher dimensions by interpreting the Rényi entropies from partition functions of replicated space-time manifolds [41, 42, 43, 27, 44]. It is in the work of EE computation in SU($N$) $J_{1}$-$J_{2}$ spin model [22], that we find although the existing EE computation algorithms can efficiently access Rényi entropy of SU($N$) spin models at small $N$ [45, 46, 30, 37, 47], the simulations become cumbersome when $N$ is large, with larger auto-correlation time and noisier data compared with their SU(2) or SU(3) cousins. The basic problem is that the configuration space in the Monte Carlo sampling process becomes exponentially large as $N$ increases, i.e., each site has $N$ local degrees of freedom, such that traditional stochastic series expansion (SSE) operator loop update [17, 48] has to deal with too many types ($N$ colors) of loops, making it exponentially hard to visit all the configurations, compared with, for example, the SU(2) case. Due to this difficulty, one cannot even compute the local observables, such as the magnetization and VBS order parameters accurately at large-$N$, let alone non-local observables, such as EE where one needs to replicate the partition function, and whose evaluation $\langle e^{-S^{(2)}}\rangle$ where $S^{(2)}$ is the 2nd order Rényi EE could render exponential explosion of its variance [49, 50].
@@
Apparently, a new QMC updated scheme is needed to reduce the complexity introduced by the multi-color loops in the SU($N$) or loop-gap models, and recently, there appears to be such a scheme that serves the purpose [51]. In the resummation-based QMC method, as schematically shown in Fig. 1, Desai and Pujari introduced the finite-$T$ operator-loop updates which directly handle uncolored loops without any reference to the underlying spin states and, therefore, manage to re-sum all the contributions from numerous colored loops into one single uncolored loop. This method works well in medium-to-large $N$ cases and, more importantly, inspires us to develop a more efficient sampling algorithm for the EE computation in the SU($N$) quantum spin models, dubbed ResumEE – resummation-based quantum Monte for entanglement entropy computation.
@@
Figure 1: Concept sketch for the resummation QMC. In this example, $N^{n_{l}=4}$ colored spin configurations with the same loop structure (panel (a)) for a general SU($N$) model are resumed into a compact uncolored loop configuration (panel (b)). In the coordinate, $r$ stands for the spatial direction of the lattice, and the $\tau$ is the imaginary time direction in the partition function.
@@
The ResumEE algorithm, therefore, not only demonstrates its superior performance compared with the previous EE computation schemes but offers a sensitive non-local probe that could distinguish the fundamental information of the transition that is otherwise very difficult for conventional probes. We, therefore, believe our ResumEE makes an essential step towards solving the critical problem of precisely evaluating the quantum entanglement in many-body systems and is of unusual intrinsic interest to a broad audience from condensed matter to quantum information and computation.
@@
where $\{b_{m},\mu_{m}\}$ identifies the $m$-th operator at bond $b_{m}$ in the operator string with operator type $\mu_{m}$ (diagonal or off-diagonal) and $|\sigma\rangle$ is the basis state the operators $H_{b_{m},\mu_{m}}$ are operated upon (usually it is the product state of the on-site $|S^{z}_{i}\rangle$). For the SU($N$) spin model, typical configurations of Eq. (1), are shown in Fig. 1 (a), one can image the operator loops as closely packed colored loops [55]. And since there are $N$ colors, the total number of configurations is exponentially large, i.e., $N^{n_{l}}$, where $n_{l}$ is the number of loops.
The resummation QMC [51] sums over the spin or color indices of these closely packed loops without changing loop connections along the operator string, such that colored configurations that share the same loop connection, as shown in Fig. 1 (a), are compressed into one uncolored loop configuration in Fig. 1 (b). The simplified partition function is now
@@
where one only needs to record the operator position $b_{m}$. In our SU($N$) Hamiltonian, $h_{b_{m}}=-\frac{J}{N}$ is the spin-symmetric matrix element, and $n_{l}$ is the number of loops in a given configuration. We emphasize that the uncolored nature of the loops emerges in the presence of SU($N$) symmetry, and the diagonal and off-diagonal operators in Eq. (1) contribute the same element in Eq. (2). Another interesting observation is that the symmetry index $N$ is now a parameter that can be tuned continuously instead of only the integer in its original form.
@@
The ResumEE algorithm.—In the ResumEE, the situation is different. Here, one can effectively sample $N^{n_{l}}$ spin configurations with each loop configuration. We can, therefore, sample a small but finite value of $\langle P_{A\rightarrow\emptyset}\rangle$ and $\langle P_{\emptyset\rightarrow A}\rangle$ each time instead of waiting for a 1 to appear indefinitely. To be more specific, the two probabilities $\langle P_{A\rightarrow\emptyset}\rangle$ and $\langle P_{\emptyset\rightarrow A}\rangle$ can be evaluated as the ratios of numbers of possible spin configurations under different boundary conditions. They can be measured with the reduction of the number of loops when one enforces the additional boundary condition, as the number of loops determines the number of possible spin configurations. That is,
@@
Fig. 2 (c) explicitly shows an example of measuring $P_{\emptyset\rightarrow A}$ according to Eq. (6). Before gluing the loops into the $C\in A$ manifold, one has $n_{1}+n_{2}=7$. After applying the additional boundary condition, one has $n_{glued}=5$. Therefore, one sample a $P_{\emptyset\rightarrow A}=N^{-2}$ with this loop configuration, as opposed to sample a $1$ every $N^{2}$ times if one were to measure with the original method with spin configurations. $P_{A\rightarrow\emptyset}$ can be measured in a similar manner, with the only difference being that the additional boundary condition is cast on the top and bottom of one of the replicas.
In practice, one only needs to count the number of distinct groups reduced at the boundary of imaginary time during such a gluing process, as the loops in the middle of time propagation will appear in both cases and cancel out in the ratio of partition functions, e.g., loops 2, 4 and 7 in Fig. 2 (c). Since the number of spin configurations measured in the resummation algorithm increases exponentially with system size, the ResumEE succeeded in mitigating the exponential decay in EE, or the exponential explosion of its variance [49, 50, 56], as it happened in the previous methods [43, 27, 44]. We now turn to the results to demonstrate such superior performance.
@@
Phase transition in the SU($N$) Heisenberg model with continuously varying $N$.— The more distinctive feature of ResumEE is its ability to compute EE efficiently at non-integer $N$ values. Then, we turn to the quantum phase transition in the SU($N$) model as a function of continuously varying $N$. It is known that this model has a Néel ordered ground state at SU($N\leq 4$) while a VBS ground state at SU($N\geq 5$[3, 14, 57]. Later, Ref. [16] with projection QMC in the VBS basis suggested a continuous phase transition at $N_{c}\approx 4.57(5)$ separating the Néel phase with onsite spin rotational symmetry breaking and the VBS phase with $Z_{4}$ lattice symmetry breaking. This finding is interesting because the transition, if continuous, shares the flavors of deconfined quantum criticality (DQC) in that two spontaneously symmetry-breaking phases meet at a single critical point with one tuning parameter. Although we have to acknowledge that the present case is different from the DQC in that the symmetry of the Hamiltonian changes when $N$ changes, the Néel-VBS transition in the present case could nevertheless offer new ingredients of novel quantum phase transitions.
@@
Here, we choose $\mathbf{q_{1}}=(2\pi/L,0)$ and $\mathbf{q_{2}}=(4\pi/L,0)$ for better data quality and note that we only keep $\chi_{\mathrm{VBS}}$ data up to $L=32$ because measurements of $C_{V}$ become too noisy at large $N$ and $L$. One sees from Fig. 4 (a) and (b) that both the dimensionless quantity $R_{N}$ and $\xi_{\mathrm{VBS}}/L$ have crossing points for various system sizes. A power law fitting of crossings for adjacent system sizes eventually points to a single point, $N_{c}=4.43(2)$ at the thermodynamic limit, $01/L=$ 0, as shown in the inset of Fig. 4 (b). A critical point near $N_{c}\approx 4.43$ can also be seen from the scaling of order parameters as a function of system sizes. One can see from Fig. 4 (c) and (d) that at $N=4.44\approx N_{c}$, both $\chi_{V}(\mathbf{X})/\beta$ and $S_{N}(\mathbf{\Gamma})$ vanish in power law of system sizes, indicating a phase transition, although we note the exponents in the power law are different, and it means that the in this case, $\eta_{N}\neq\eta_{V}$. Although there is no lucid indication for a first-order transition, it is worth suspecting a weakly-first-order transition analogous to several DQC scenarios at small $N$[22, 30, 20, 19]. Distinguishing a weakly first-order transition from a continuous one is a formidable task. It can not definitely be answered with the measurement of local observables, even though our $L$ is already larger than the previous work.
Figure 5: Scaling of entanglement entropy at different $N$ in the SU($N$) model. Panel (a) shows $S^{2}_{A}(l_{A})$ obtained from the ResumEE method with smooth boundary at the transition point $N_{c}=4.43$ determined from the Binder ratio crossings in Fig. 4. A clear sub-leading logarithmic correction is seen with the coefficient $b=1.76(4)$ (panel (b)). We also show the EE for SU(2.4), SU(2.8), SU(3), SU(3.2) and SU(4), these systems are well inside the Néel phase with $N_{G}=2(N-1)$ Goldstone modes, such that their sub-leading logarithmic corrections yields $b=\frac{N_{G}}{2}=N-1$, with $b=1.35(6),1.68(8),1.9(1),2.15(8),2.88(9)$ for the SU(2.4), SU(2.8), SU(3), SU(3.2) and SU(4) cases in panel (b). Note that for SU($4$), we consider the $J_{1}-J_{2}$ Hamiltonian (dashed line in (a) and (b)) to enhance the Néel order without changing universal information. Panel (c) summarizes the obtained $b$ with respect to $N$, and the black line is $b=N-1$ according to the Goldstone modes inside the SU(N) symmetry-breaking phase. The red dot is the $b$ for SU(4.43), whose finite value suggests the coexistence of the Néel state with the VBS state at the transition point. Note that $b$ at $N=2$ adopted from Ref. [60] is also presented in panel (c).
@@
In order to strengthen our argument, we also investigate the EE scaling inside the Néel phase, $N<N_{c}$ using ResumEE. Note that as $N$ approaches $N_{c}$, the finite-size effect in EE scaling becomes stronger. We introduce an extra next-nearest neighbor $J_{2}$ term, $-J_{2}\sum_{\langle\langle ij\rangle\rangle}\Pi_{ij}$ to Eq. (7) at SU(4) case to reduce the finite-size effect, where $\Pi_{ij}=\frac{1}{N}\sum_{\alpha,\beta}|\beta_{i}\alpha_{j}\rangle\langle\alpha_{i}\beta_{j}|$ is the SU($N$) generalization of the next-nearest-neighbor ferromagnetic interaction and, therefore, enhances the Néel order while keeping universal information, such as the number of Goldstone modes, $N_{G}$, intact. In SU(4) case (shown as a dashed line in Fig. 5), we set $J_{2}=3.5$ [46]. As shown in Fig. 5 (b), staying deep inside the Néel phase, we successfully detect the correct number of Goldstone modes $N_{G}=2(N-1)$ at SU(3) and SU(4) from EE scaling with the log coefficient, $b=1.9(1)$ and $2.88(9)$ respectively. More interestingly, our ResumEE results find the consistency of $N_{G}=2(N-1)$ at non-integer $N$ cases, $N=2.4,2.8,3.2$ with $b=1.35(6),1.68(8),2.15(8)$ respectively, as shown in Fig. 5 (c). This is another novel result from ResumEE that the CFT prediction of the log-correction can be analytically continued to non-integer $N$s, or in other words, the number of Goldstone modes $N_{G}$ can also be fractionalized. Therefore, with the observation inside the Néel ordered phase, the finite log-correction at $N_{c}$ serves as a clue of continuous symmetry-breaking right at this point, eventually suggesting a weakly-first-order phase transition [62, 63].
@@
Discussion.—The precise computation of the EE for quantum many-body systems is the long-standing obstacle that blockades our full-fledged exploration of the fundamental entanglement and CFT information of quantum matter. In the continuous efforts over the past two decades, various estimators of EE all suffered from the exponential explosion of its variances when the entanglement region just starts to scale [26, 52, 53, 54, 42, 43, 27, 44, 50, 49]. In this work, based on the recently developed resummation-based quantum Monte Carlo method for the SU($N$) spin and loop-gas models [51], we develop a new algorithm – ResumEE – manage to mitigate this long-standing obstacle. Our ResumEE exponentially speeds up the computation of the exponentially small value $\langle e^{-S^{(2)}_{A}}\rangle$ such that the EE for a generic quantum SU($N$) spin models at any continuous $N$ values can be readily computed.
We demonstrate the superior performance of our algorithm over the previous estimators of $S^{(2)}_{A}$ on 1D and 2D SU($2$) Heisenberg spin systems. More importantly, we use it to detect the entanglement scaling data of the Néel-VBS transition on 2D SU($N$) Heisenberg model on a square lattice with $continuously$ varying $N$. Our EE results reveal that although previous work suggested there is a continuous transition at $N_{c}\approx 4.57(5)$ [16], the transition deviates from a CFT description with a substantial log-correction to the area law scaling in EE when the entanglement region has a smooth boundary. Such log-correction may come from the domains that exhibit SU($N$) continuous symmetry-breaking with gapless Goldstone modes, which coexist with the VBS domains at their weakly-first-order transition. Our local order parameter measurements convey the same message of phase coexistence of the Néel and VBS states once pushed to much larger system sizes. Our ResumEE also reveals a non-trivial phenomenon: the log-corrections of EE inside the Néel phase at non-integer $N$ also follow the CFT prediction, leading to the fractionalization of Goldstone modes. The ResumEE algorithm, therefore, makes an essential step towards extracting the quantum entanglement data and understanding the novel quantum criticality and encourages more works in many-body systems with SU($N$) symmetry at non-integer $N$ values.